# Preempting Fermion Sign Problem: Unveiling Quantum Criticality through Nonequilibrium Dynamics

### v1

@@ Abstract
where $\alpha=1$, $2$, $3$ is the flavor index of fermion, $U$ is repulsive Hubbard interaction strength, and $t_{ij}=te^{i\theta_{ij}}$, in which $t=1$ is set as energy unit. As illustrated in Fig. 3a, the magnetic flux in each plaquette is $\sum_{\square}\theta_{ij}=(-1)^{i_{x}+i_{y}}\phi$. For non-zero $\phi$, the energy dispersion of Eq. (4) in the non-interacting limit features two Dirac points located at momenta (±$(\pm\frac{\pi}{2},\pm\frac{\pi}{2})$, as shown in SM [55]. We fix our simulation at half filling such that the Fermi level is located at Dirac points. The model of Eq. (2) respects $\rm SU(3)\times Z_{2}$ symmetry, where $\rm SU(3)$ is the rotation symmetry in flavor space and $\rm Z_{2}$ is the sublattice symmetry which exchanges the two sublattices on the staggered-flux square lattice with each unit cell containing two sites.
The $\rm SU(N)$ symmetry plays overarching roles in modern physics. For instance, the $\rm SU(3)$ symmetry lays the foundation for strong interaction between quarks [3]. Recently, fermionic interacting models with $\rm SU(N)$ symmetry have been realized in optical lattice [61, 62]. Extensive theoretical and numerical efforts, including sign-free QMC simulation, are devoted to understanding the exotic phenomena arising from the interplay between strong interaction, Dirac fermions and multi-flavor physics [63, 64, 65, 66, 67]. However, despite its fundamental importance, the unbiased numerical study on $\rm SU(N)$-symmetric model with repulsive Hubbard interaction for odd $N$ is scarce, largely due to the presence of notorious sign problem in QMC simulation. Hence, we implement our approach to preempt sign problem and unravel the quantum phases and exotic QCP in Eq. (4), which is a minimal interacting model featuring $\rm SU(N)$ symmetry with odd $N$.
@@
In the $\lambda_{8}$-AFM ordered phase, the $\rm Z_{2}$ sublattice symmetry is broken and the $\rm SU(3)$ symmetry is broken into $\rm SU(2)\times U(1)$ (See SM Sec. IV for details [55]), as illustrated in Fig. 3a. Accordingly, the order parameter manifold is $\rm\frac{SU(3)\times Z_{2}}{SU(2)\times U(1)}$, which is remarkably different from the universality classes of the Gross-Neveu QCP studied previously. Thus, the QCP separating DSM and $\lambda_{8}$-AFM phase is a novel Dirac QCP.
@@
To reveal the critical properties, we perform short-time QMC simulation with fully ordered initial state by fixing $\phi=0.075\pi$ and $\tau L^{-z}=0.25$ where $z=1$. The results of correlation length ratio $R_{\lambda_{8}\text{-AFM}}$ for $\lambda_{8}$-AFM order (defined in SM Sec. IV [55]) are shown in Fig. 3b and the crossing points dictate the phase transition point from DSM to $\lambda_{8}$-AFM phase occurring at $U_{c}=1.10(5)$. We also compute the correlation-length ratios for other types of AFM order instead of $\lambda_{8}$, further verifying that only $\lambda_{8}$-AFM long-range order is present in the ground state [55]. With varying magnetic flux $\phi$, we implement a similar procedure and arrive at the ground-state phase diagram of the Eq. (4), as shown in Fig. 3a, in which $U_{c}$ increases with $\phi$.
@@
After accessing the phase diagram, we investigate the critical properties. The data collapse analysis of $R_{\lambda_{8}\text{-AFM}}$ versus $(U-U_{c})L^{1/\nu}$ in the regime close to QCP with fixed $\phi=0.075$ gives $1/\nu=0.68(5)$, as shown in Fig. 3c. Moreover, at $U_{c}$, the short-time dynamics of the structure factor $S_{\lambda_{8}\text{-AFM}}$ and the fermion correlation function $G_{\rm f}$ (defined in SM Sec. IV [55]) yield $\eta_{\phi}=0.55(5)$ and $\eta_{\psi}=0.15(3)$, as presented in Fig. 3d and 3e, respectively. To confirm the universality of this phase transition, we also evaluate the critical exponents for other values of $\phi$, as presented in SM [55], which are consistent with each other within the error bar. Remarkably, the critical exponents found here are different from those of the Gross-Neveu universality classes [69, 70] (for comparison, see SM [55]), confirming that this phase transition belongs to a new universality class.
Concluding remarks — In summary, we establish an innovative theoretical framework to unravel critical properties in quantum many-body systems. The approach proposed in our work enables preempting the notorious sign problem in QMC by leveraging the short-stage imaginary-time relaxation dynamics. We demonstrate the accuracy and efficiency of the approach by systematically investigating several strongly interacting models by sign-problematic QMC. The comparison of our approach with conventional QMC shows that the sign problem is enormously reduced in our approach, thus largely reducing the computational time and enabling the accessing the numerically accurate critical exponents. For example, for the $\rm SU(3)$ Hubbard model with $L=10$, the estimated computational time in equilibrium QMC is millions of times that in our approach [55]. Consequently, our work paves a novel avenue to studying exotic QCP in quantum many-body systems, which are not amendable to the unbiased numerical approach previously.
@@
More intriguingly, employing the newly developed approach, for the first time we reveal the ground-state properties of the staggered-flux $\rm SU(3)$ repulsive Hubbard model, and reveal a novel Dirac QCP different from previously known Gross-Neveu transitions. Our numerical studies motivate systematic theoretical analysis of the QCP in $\rm SU(3)$ Dirac fermions by field-theory approach. Moreover, the critical properties are potentially detectable in the synthetic quantum simulators such as optical lattice of cold atoms [61, 62]. Our results on the $\rm SU(3)$ Hubbard model provide theoretical guidance for the future experimental exploration of exotic physics in strongly correlated $\rm SU(N)$ Dirac fermions.
The basic procedure in our approach can be straightforwardly generalized to other fermionic QCP, including metallic QCP involving Fermi surface, for which the critical properties attract considerable attentions, but remain largely elusive. Additionally, we believe that short-time nonequilibrium QMC is also applicable in sign problematic bosonic models. We should mention that in some specific cases, although sign problem is alleviated, it remains difficult to access the critical properties with large system size in our current simulation. For example, when staggered magnetic flux is large in $\rm SU(3)$ Hubbard model, even at short-time stage the sign problem is too severe to allow us reach the results with high accuracy close to the QCP. Nonetheless, bestowing the flexibility of the framework, it is feasible to further ameliorate the efficiency of our approach, for instance through improvement of initial state and optimization of H-S channel [32, 31], which is an interesting direction in future. We believe that our work promises a pathway towards a general strategy of studying the QCP in strongly correlated systems by unbiased numerical approach.
@@
Acknowledgments — We acknowledge helpful discussions with Yiwen Pan. Yin-Kai Yu, Zhi-Xuan Li and Shuai Yin are supported by the National Natural Science Foundation of China (Grants No. 12222515 and No. 12075324). Zi-Xiang Li is supported by the NSFC under Grant No. 12347107. Shuai Yin is also supported by the Science and Technology Projects in Guangdong Province (Grants No. 2021QN02X561).
@@
Preempting Fermion Sign Problem: Unveiling Quantum Criticality through Nonequilibrium Dynamics
I Details of nonequilibrium imaginary-time dynamics in PQMC simulation
I.1 Imaginary-time relaxation dynamics simulated by PQMC
We focus on the imaginary-time relaxation dynamics from a fully ordered or Dirac semimetal initial state |$\ket{\psi_{0}}$. The initial state is prepared by solving $H_{0}\ket{\psi_{0}}=E_{0}\ket{\psi_{0}}$, where $H_{0}$ is the initial Hamiltonian and $E_{0}$ is the ground state energy. With these initial states, the evolution of the observable $O(\tau)$ is given by
@@ I.1 Imaginary-time relaxation dynamics simulated by PQMC
The imaginary-time relaxation dynamics can be simulated via the projector quantum Monte Carlo (PQMC) [51]. In conventional PQMC studies, a sufficiently large $\tau$ (usually $\tau$ should be several times as large as $L$) is needed to ensure that the ground state is obtained. Then the physical qualities are calculated in the ground state. In contrast, in our work, we focus on the short-time stage of the imaginary-time relaxation process, and $\tau$ does not need to be very large compared with $L$.
I.2 Hubbard-Stratonovich transformation
In the PQMC simulations, the interaction terms in the form of four-fermion operators should be decoupled via Hubbard-Stratonovich (HS) transformation. We at first use the Trotter decomposition to discretize the imaginary-time evolution operator into ${M=\tau/\Delta\tau}$ (where ${M}$ is an integer) time slices, i.e.,
@@ I.2 Hubbard-Stratonovich transformation
where $\cosh\lambda=\mathrm{e}^{\frac{\Delta\tau U}{2}}$. For $U>0$ (Hubbard repulsive interaction), the sign problem arises for all channels. We here choose the $\sigma_{x}$ channel to mitigate the sign problem [52].
@@
I.3 Sign problem in PQMC
@@
$\braket{O}=\frac{\sum_{\mathrm{c}}w(\mathrm{c})O(\mathrm{c})}{\sum_{\mathrm{c}}w(\mathrm{c})}=\frac{\sum_{\mathrm{c}}|w(\mathrm{c})|\mathrm{sign}(\mathrm{c})O(\mathrm{c})/\sum_{\mathrm{c}}|w(\mathrm{c})|}{\sum_{\mathrm{c}}|w(\mathrm{c})|\mathrm{sign}(\mathrm{c})/\sum_{\mathrm{c}}|w(\mathrm{c})|}=\frac{\braket{O}_{|w|}}{\braket{\mathrm{sign}}_{|w|}}.$ (S6)
@@
II More details for the single-Dirac-fermion Hubbard model
II.1 Sign problem behaviors

### v2

@@ RESULTS
where $\alpha=1$, $2$, 33 is the flavor index of fermion, $U$ is repulsive Hubbard interaction strength, and $t_{ij}=te^{i\theta_{ij}}$, in which $t=1$ is set as energy unit. As illustrated in Fig. 3a, the magnetic flux in each plaquette is $\sum_{\square}\theta_{ij}=(-1)^{i_{x}+i_{y}}\phi$. For $\phi=0$ cases, recent constrained-path QMC and finite-temperature determinant QMC studies have investigated the model at 1/3 filling [64, 65]. For non-zero $\phi$, the energy dispersion of Eq. (4) in the non-interacting limit features two Dirac points located at momenta $(\pm\frac{\pi}{2},\pm\frac{\pi}{2})$, as shown in the inset of Fig. 3a. We fix our simulation at half filling such that the Fermi level is located at the Dirac points. The model of Eq. (4) respects $\rm SU(3)\times Z_{2}$ symmetry, where $\rm SU(3)$ is the rotation symmetry in flavor space and $\rm Z_{2}$ is the sublattice symmetry which exchanges the two sublattices on the staggered-flux square lattice with each unit cell containing two sites.
The $\rm SU(N)$ symmetry plays overarching roles in modern physics. For instance, the $\rm SU(3)$ symmetry lays the foundation for strong interaction between quarks [3]. Recently, interacting fermionic models with $\rm SU(N)$ symmetry have been realized in optical lattice [66, 67]. Extensive theoretical and numerical efforts, including QMC simulations, are devoted to understanding the exotic phenomena arising from the interplay between strong interaction, Dirac fermions and multi-flavor physics [68, 69, 70, 71, 72]. However, despite its fundamental importance, unbiased numerical studies on $\rm SU(N)$-symmetric model with repulsive interaction for odd $N$ are scarce, primarily due to the notorious sign problem. Hence, we implement our approach to preempt the sign problem and unravel the quantum phases and exotic QCP in Eq. (4), a paradigmatic interacting model for odd NN $\rm SU(N)$ symmetry.
@@
In the $\lambda_{8}$-AFM ordered phase, the $\rm Z_{2}$ sublattice symmetry is broken and the $\rm SU(3)$ symmetry is broken into $\rm SU(2)\times U(1)$ (See SM Sec. S4.4 for details [53]), as illustrated in Fig. 3a. This order parameter transforms under the adjoint representation of SU(3), spanning the vacuum manifold $\frac{\text{SU(3)}\times\text{Z}_{2}}{\text{SU(2)}\times\text{U(1)}}$. Such symmetry breaking defines a new unconventional Gross-Neveu QCP, where the order parameter symmetry is SU(3), in contrast to the O($N$) order parameters of the conventional Gross-Neveu universality classes such as chiral Ising, chiral XY, and chiral Heisenberg.
@@
To reveal the critical properties, we perform short-time QMC simulation starting from a fully ordered state at $\phi=0.075\pi$ and $\tau L^{-z}=0.25$ ($z=1$). The results of correlation length ratio $R_{\lambda_{8}\text{-AFM}}$ for $\lambda_{8}$-AFM order (defined in SM Sec. S4 [53]) are shown in Fig. 3b and the crossing points dictate the phase transition point from DSM to $\lambda_{8}$-AFM phase occurring at $U_{c}=1.10(5)$. We also compute the correlation-length ratios for other types of AFM order instead of $\lambda_{8}$, confirming that only $\lambda_{8}$-AFM long-range order is present in the ground state [53]. With varying magnetic flux $\phi$, we apply the same procedure to map out the ground-state phase diagram of the Eq. (4), as shown in Fig. 3a, in which $U_{c}$ increases with $\phi$.
@@
After accessing the phase diagram, we investigate the critical properties. The data collapse analysis of $R_{\lambda_{8}\text{-AFM}}$ versus $(U-U_{c})L^{1/\nu}$ in the regime close to QCP with fixed $\phi=0.075$ gives $1/\nu=0.68(5)$, as shown in Fig. 3c. Moreover, at $U_{c}$, the short-imaginary-time dynamics of the structure factor $S_{\lambda_{8}\text{-AFM}}$ and the fermion correlation function $G_{\rm f}$ (defined in SM Sec. S4 [53]) yield $\eta_{\phi}=0.55(5)$ and $\eta_{\psi}=0.15(3)$, as presented in Fig. 3d and 3e, respectively. To confirm the universality of this phase transition, we also evaluate the critical exponents for other values of $\phi$, as presented in SM [53], and they are mutually consistent within error bars. Remarkably, the critical exponents found here are numerically distinct from those of any conventional Gross-Neveu universality class [74, 75] (for comparison, see SM [53]), providing conclusive evidence that this transition belongs to a new universality class.
DISCUSSION
In summary, we introduce an innovative theoretical framework unraveling ground-state properties, with a particular focus on quantum criticality, in quantum many-body systems. Our method effectively circumvents the notorious sign problem in QMC by leveraging the short-stage imaginary-time relaxation dynamics. We provide compelling evidence of its accuracy and efficiency across several strongly interacting models, showcasing a dramatic reduction in computational cost compared to conventional equilibrium QMC. For example, the estimate time for the $\rm SU(3)$ Hubbard model with $L=10$ is billions of times shorter with our approach than with equilibrium PQMC [53]. Furthermore, the inherent flexibility in choosing initial states, coupled with a robust scaling relation, provides a crucial benchmark for ensuring accuracy within our framework. Besides, this approach can distinguish between first-order and continuous phase transitions [53]. This work, therefore, establishes a powerful strategy for the unbiased numerical study of the exotic quantum phases and quantum phase transitions in a broad class of previously computationally inaccessible systems.
@@ DISCUSSION
More intriguingly, we leverage our newly developed approach to make a pioneering investigation into the $\rm SU(3)$ repulsive Hubbard model hosting Dirac fermions. It represents the first time that the ground-state phase diagram and the QCP in the $\rm SU(N)$ repulsive Hubbard model with odd $N$ have been unveiled through unbiased numerical approaches. We identify the first example of Gross-Neveu QCP that fundamentally differs from the conventional ones typically confined to the chiral O($N$) universality classes [53]. Moreover, the critical properties are potentially detectable in the synthetic quantum simulators such as optical lattice of cold atoms [66, 67]. Hence, our findings on the $\rm SU(3)$ Hubbard model not only establish a significant novel direction in modern statistics physics for exploring the fermionic QCP but also offer critical theoretical guidance for the forthcoming experimental exploration of exotic physics in strongly correlated $\rm SU(N)$ fermionic models.
Our innovative approach presents a powerfully generalizable framework, extending readily to other fermionic QCP, including the notoriously challenging metallic QCP involving Fermi surfaces, whose critical properties are of immense theoretical and experimental interest but have remained largely inscrutable. Moreover, the short-time nonequilibrium QMC is also applicable to sign-problematic bosonic models, opening entirely new avenues. We carefully delineate the boundaries of our method: it is not a fully generalized solution to the sign problem. Arbitrarily short evolution times retain excessive memory of the initial state, preventing evolution into the desired nonequilibrium scaling regime (see SM Sec. S1 [53] for detailed discussions). This means a mild sign problem ensuring accurate results within the short-time nonequilibrium scaling regime cannot be universally guaranteed for every model. For instance, in the SU(3) Hubbard model with large staggered magnetic flux, the sign problem remains too severe for highly accurate results near the QCP, even in the short-time stage. Despite these limitations, the framework’s flexibility allows for future efficiency enhancements, such as through improvement of initial state and optimization of H-S channel [32, 31]. This work is not merely an incremental improvement; it delivers a novel pathway toward establishing a general and unbiased numerical strategy for unraveling the mysteries of the ground-state phase diagram and QCP in strongly correlated systems, a long-sought goal in the field.
Acknowledgments: We acknowledge helpful discussions with Yiwen Pan, Chengfeng Cai, Shao-Hang Shi, and Chang-Yu Shen.
@@
Funding Statement: Yin-Kai Yu, Zhi-Xuan Li, and Shuai Yin are supported by the National Natural Science Foundation of China (Grants No. 12222515 and 12075324). Zi-Xiang Li is supported by the National Natural Science Foundation of China under Grant No. 12347107. Shuai Yin is also supported by the Science and Technology Projects in Guangdong Province (Grant No. 2021QN02X561) and in Guangzhou City (Grant No. 2025A04J5408).
Author Contributions Statement: Shuai Yin and Zi-Xiang Li conceived and supervised the project. Yin-Kai Yu and Zhi-Xuan Li carried out the numerical calculations and data analysis. All authors participated in discussions and contributed to writing the manuscript.
Competing Interests Statement: The authors declare they have no competing interest.
Data & Materials Availability Statement: All raw data underlying the results presented in the main text and Supplementary Materials are archived in the Dryad Digital Repository (DOI: 10.5061/dryad.dncjsxmch) under the CC0 Public Domain Dedication.
@@
Preempting Fermion Sign Problem: Unveiling Quantum Criticality through Nonequilibrium Dynamics in Imaginary Time
S1 Details of nonequilibrium imaginary-time dynamics in PQMC simulation
S1.1 Imaginary-time relaxation dynamics simulated by PQMC
[…]
@@
The imaginary-time relaxation dynamics can be simulated via the projector quantum Monte Carlo (PQMC) [54, 55, 51]. In conventional PQMC studies, a sufficiently large $\tau$ (usually $\tau$ should be several times as large as $L$) is needed to ensure that the ground state is obtained. Then the physical qualities are calculated in the ground state.
In contrast, in our work, we focus on the short-time stage of the imaginary-time relaxation process, where $\tau$ does not need to be large compared to $L$. In this regime, the system remains in a nonequilibrium state rather than the ground state This stage is usually discarded in conventional PQMC studies. In our previous work, we systematically and carefully investigated the imaginary-time nonequilibrium dynamics of Dirac fermion systems [50], and found that although this setup goes beyond the conventional PQMC framework, choosing a small $\tau$ only introduces some dependence on the initial state, while the ground-state information, particularly universal quantum criticality, still manifests in the nonequilibrium dynamics. With a properly chosen initial state, one can fix $\tau/L^{z}$ to a small value and still observe quantum phase transitions via standard finite-size scaling analysis, even in this nonequilibrium regime.
A similar practice has also appeared earlier in the context of finite-temperature QMC (FTQMC). In FTQMC, the expectation value of an observable is given by $\braket{O}=\frac{1}{Z}\text{Tr}[\mathrm{e}^{-\beta(H-\mu N)}O]$, with $Z=\text{Tr}[\mathrm{e}^{-\beta(H-\mu N)}]$ [5, 6, 76]. Here, $\beta=1/k_{\text{B}}T$ is the inverse temperature, and it corresponds to the imaginary-time extent in the Monte Carlo simulation. Unlike zero-temperature PQMC, FTQMC involves both quantum and thermal fluctuations. When FTQMC is used to study zero-temperature quantum phase transitions, it is common to fix $\beta/L^{z}$ to a finite value (typically $\beta=L^{z}$) to obtain temperature-independent finite-size scaling. There are also studies using smaller values, such as $\beta=\frac{1}{4}L^{z}$ [77].
Although both $\beta$ and $\tau$ represent the imaginary-time length in QMC simulations, they carry distinct physical meanings. A shorter $\beta$ in FTQMC corresponds to a higher-temperature thermal equilibrium state, and it is widely accepted that $\beta/L^{z}$ can be used as a dimensionless scaling factor. In contrast, a shorter $\tau$ in zero-temperature PQMC corresponds to a nonequilibrium state and has rarely been considered outside the context of imaginary-time nonequilibrium dynamics [49, 78, 50]. Here, based on our understanding of imaginary-time nonequilibrium dynamics, we leverage this idea to address the fermion sign problem in zero-temperature PQMC simulation. Moreover, compared to conventional PQMC and FTQMC approaches, our framework further alleviates the sign problem by flexibly choosing appropriate initial states, while the consistency of results obtained from different initial states provides a useful means of self-consistency check.
S1.2 Hubbard-Stratonovich transformation
In the PQMC simulations, the interaction terms in the form of four-fermion operators should be decoupled via Hubbard-Stratonovich (HS) transformation. The implementation of PQMC begins with performing the Trotter decomposition to discretize the imaginary-time evolution operator into $M=\tau/\Delta\tau$} (where M{M} is an integer) time slices, i.e.,
@@
where $\cosh\lambda=\mathrm{e}^{\frac{\Delta\tau U}{2}}$. For $U>0$ (Hubbard repulsive interaction), the sign problem arises for all channels. We here choose the $\sigma_{x}$ channel to mitigate the sign problem [56]. This choice also affects the structure of the fermion determinant. Specifically, due to the spin-flipping nature of the free-fermion hopping term in the single-Dirac-fermion Hubbard model, the $\sigma_{x}$-channel decoupling leads to an auxiliary-field Hamiltonian that only has off-diagonal blocks. As a result, the total fermion determinant can be factorized into the product of two smaller determinants: $\det(M)\propto\det(M_{\uparrow\downarrow})\det(M_{\downarrow\uparrow})$. In contrast, decoupling in the $\sigma_{z}$ channel does not permit this factorization. This results in a single determinant of twice the dimension, which typically exacerbates the sign problem.
@@
S1.3 Sign problem in PQMC
@@
$\braket{O}=\frac{\sum_{\mathrm{c}}w(\mathrm{c})O(\mathrm{c})}{\sum_{\mathrm{c}}w(\mathrm{c})}=\frac{\sum_{\mathrm{c}}|w(\mathrm{c})|\mathrm{sign}(\mathrm{c})O(\mathrm{c})/\sum_{\mathrm{c}}|w(\mathrm{c})|}{\sum_{\mathrm{c}}|w(\mathrm{c})|\mathrm{sign}(\mathrm{c})/\sum_{\mathrm{c}}|w(\mathrm{c})|}=\frac{\braket{O\,\mathrm{sign}}_{|w|}}{\braket{\mathrm{sign}}_{|w|}}.$ (S6)
@@
S1.4 Applicability and limitations of the nonequilibrium approach
The proposed nonequilibrium method offers an efficient pathway for exploring ground-state properties by circumventing the sign problem in our model and significantly lowering computational cost. This makes it highly effective for mapping the phase diagram and extracting critical properties of the quantum phase transition. However, it is important to emphasize that this is not a universal solution for all models plagued by the sign problem. Here, we discuss the applicability and limitations of our approach. The nonequilibrium approach is subject to two key limitations that constrain its applicability:
  1. 1.
    One cannot take the imaginary time arbitrarily short in order to ensure that the sign problem is mild for all models. If the evolution time is too short, the system retains too much memory of the initial state, and the scaling theory breaks down. For instance, this breakdown of scaling theory occurs when the characteristic length scale of the evolution, $\xi_{l}\propto\tau^{1/z}$, becomes comparable to the system’s ultraviolet (UV) cutoff—in our case, the lattice spacing (set to unity). Under this condition, the behavior of observables is dominated by non-universal, high-energy physics, rendering the scaling analysis invalid.
  2. 2.
    This method is applicable to the continuous quantum phase transitions. This method is applicable for studying continuous quantum phase transitions. Our approach can be applied to determine the ground-state phase diagram and to identify and locate possible QCPs. If a QCP exists, our method can accurately determine the critical exponents based on the scaling theory of continuous transitions. On the other hand, if the calculation reveals that the transition does not exhibit scaling behavior, it implies a first-order transition, for which there is currently no well-established scaling analysis framework to use. In addition, our method is not applicable deep in the ordered phase, far away from the transition point.
We begin by elaborating on the first limitation, and provide a framework for understanding when the method is effective and when it fails. In general, fixing a system size $L$ for discussion, the effectiveness of our method depends on three characteristic imaginary-time scales: (1) the minimum time $\tau_{\text{neq.}}$ at which the nonequilibrium scaling theory becomes valid; (2) the minimum time $\tau_{\text{eq.}}$ required for conventional PQMC to project onto the ground state; and (3) the maximum time $\tau_{\text{sign}}$ before the sign problem becomes too severe to obtain reliable results. Empirically, $\tau_{\text{eq.}}$ varies slightly between models but is usually around two to three times the linear system size $L$. The time $\tau_{\text{neq.}}$ required for short-time scaling to apply is typically one or two orders of magnitude smaller than $\tau_{\text{eq.}}$. In contrast, $\tau_{\text{sign}}$ can differ drastically from model to model and is the dominant factor that determines whether our approach will be effective. This leads to three possible scenarios:
  1. 1.
    $\tau_{\text{neq.}}<\tau_{\text{eq.}}<\tau_{\text{sign}}$: the sign problem is mild, and equilibrium methods already work well. Our nonequilibrium method only improves computational efficiency.
  2. 2.
    $\tau_{\text{neq.}}<\tau_{\text{sign}}<\tau_{\text{eq.}}$: equilibrium methods fail due to the sign problem, but the nonequilibrium method still works. This is the type of situation our work is focused on.
  3. 3.
    $\tau_{\text{sign}}<\tau_{\text{neq.}}<\tau_{\text{eq.}}$: the sign problem is too severe even at short times. In this case, the nonequilibrium method also fails.
Because the sign problem behaves differently across models, $\tau_{\text{sign}}$ varies accordingly. Based on this classification, we can view sign-problematic strongly correlated models as falling into three broad categories. The general argument that a QCP can be approached using sufficiently short imaginary-time simulations where the sign problem remains mild is valid only in models of type 1 and type 2. It does not hold in type 3 models, where the sign problem is already severe even at short times. Naturally, our method is not a fully general solution to the sign problem in all cases. What it does provide is a practical and effective approach for models of type 2, which were previously inaccessible due to severe sign problems in equilibrium simulations. With our method, such models become tractable.
Within the same model and parameter set, $\tau_{\text{sign}}$ also depends on the system size $L$. Typically, $\tau_{\text{sign}}$ decreases as $L$ increases. This is because the average sign scales exponentially with both the particle number—which grows with system size ($N\propto L^{d}$)—and $\tau$, as described by (S8[9], although exceptions have been noted [52, 63]. Taking this size dependence into account, we can sketch a regime diagram in the $(L,\tau)$ plane with the three regimes described above, as illustrated in Fig. S1. This diagram clearly demarcates the three previously described regimes and highlights a key advantage of the nonequilibrium method: it grants access to significantly larger system sizes than those achievable with conventional equilibrium QMC.
Figure S1: Schematic illustration of the three scenarios regarding the sign problem. The shaded region indicates where the sign problem is not severe and simulations are feasible. The green region corresponds to scenario 1 ($\tau_{\text{neq.}}<\tau_{\text{eq.}}<\tau_{\text{sign}}$), where equilibrium methods also work. The yellow region corresponds to scenario 2 ($\tau_{\text{neq.}}<\tau_{\text{sign}}<\tau_{\text{eq.}}$), where only the nonequilibrium method is feasible. The red region corresponds to scenario 3 ($\tau_{\text{sign}}<\tau_{\text{neq.}}<\tau_{\text{eq.}}$), where even the nonequilibrium approach fails. The boundaries define the maximum accessible system sizes $L_{\text{max}}^{(\text{eq.})}$ and $L_{\text{max}}^{(\text{neq.})}$ for each method.
The above classification only provides general criteria for when the nonequilibrium method is effective, while the quantitative determination of the nonequilibrium scaling regime in practical applications is illustrated in detail with procedures and examples in Sec. S2.2.
Next, we discuss another boundary case of this nonequilibrium short-time method, i.e., the first-order phase transition. In the main text, we have already demonstrated that our nonequilibrium method can determine the existence of a continuous phase transition based on scaling forms. The key criteria are: 1) For a given $\tau L^{-z}$, the crossing points of the correlation-length ratio or Binder ratio curves converge to a single point as the system size increases (dimensionless quantities exhibit scale invariance at the critical point). 2) The nonequilibrium critical relaxation processes of physical quantities such as the structure factor and fermion correlation display scaling collapse.
In fact, we are also able to identify a first-order phase transition during the short-time evolution through opposite characteristics: 1) The crossing points of the correlation-length ratio or Binder ratio curves do not converge to a single point. 2) The nonequilibrium critical relaxation processes of physical quantities such as the structure factor and fermion correlation cannot be well scaled to collapse. Additionally, 3) due to the coexistence of two phases at a first-order phase transition, the Binder ratio typically shows negative dips at the transition point, and the Monte Carlo sampling distribution of the structure factor exhibits a double peak. The emergence of negative dips in the results of Binder ratio is a hallmark of the first-order transition.
As a typical example, we demonstrate how to identify the first-order phase transition in the $q=6$ quantum Potts chain during short imaginary time evolution. The Hamiltonian is given by:
$H=-qJ\sum_{i}\sum_{m=0}^{q-1}P_{i}^{(m)}P_{i+1}^{(m)}\;-\;h\sum_{i}\sum_{m\neq n}|m\rangle_{i}\langle n|\,$, (S9)
where $P_{i}^{(m)}\equiv|m\rangle_{i}\langle m|$ is the projection operator on site $i$, and the tuning parameter is $g\equiv h/J$. For $g<1$, the ground state is in the ferromagnetic phase, and for $g>1$, it is in the paramagnetic phase. For the case of $q=6$, a first-order phase transition occurs at $g_{c}=1$. We use the time-evolving block decimation (TEBD) method to simulate the system’s imaginary time evolution starting from an ordered ferromagnetic initial state (e.g., all sites choosed into $m=0$ state) and observe the order parameter $M$ and Binder ratio $R$:
$M\equiv\frac{1}{L}\sum_{i=1}^{L}s_{i}\,,\qquad R\equiv 1-\frac{\langle M^{4}\rangle}{3\,\langle M^{2}\rangle^{2}}\,$, (S10)
where $s_{i}\equiv\frac{q}{q-1}\left(P_{i}^{(0)}-\frac{1}{q}\right)$. Using the same procedure as in the main text, we take a short imaginary time $\tau/L=0.3$. The variation of Binder ratio $R$ with $g$ is shown in Fig. S2a. The curves for different system sizes do not intersect at a single point, and there are distinct negative dips, which are characteristic of a first-order phase transition and can appear with short imaginary time evolution at $\tau/L=0.3$. Fig. S2b shows the imaginary time relaxation of the order parameter $M$ starting from the ordered phase at $g=1$, which decays exponentially with $\tau$. Fig. S2c demonstrates the failure of scaling collapse for $M$, using a scaling exponent $a=1.0$ as an example. Adjusting $a$ does not allow these curves to collapse or partially collapse. If we fix $\tau/L$ and perform scaling fitting according to the form $M\propto L^{-a}$, the results for $a$ are shown in Fig. S2d, and they do not converge as $\tau/L$ increases. These results, which violate scaling forms, clearly exclude the possibility of a continuous phase transition here. This example illustrates that a first-order phase transition can be distinguished from a continuous phase transition even during short imaginary time evolution. However, for such cases, although the rough phase diagram can still be determined, the precise location of the transition point may not be accurately obtained within the short imaginary-time regime since the current scaling form no longer holds, and may instead require extrapolation from finite-time data.
Figure S2: First-order phase transition characteristics in the $q=6$ quantum Potts chain observed during short imaginary time evolution. a, Binder ratio versus $g$ for different system sizes, where the curves do not intersect at a single point, and negative dips are observed. The emergence of negative dips in the Binder ratios is a hallmark of first-order transition. b, Imaginary time relaxation of the order parameter $M$ starting from the ordered phase at $g=1$. c, Failure of scaling collapse for b, with $a=1.0$ as an example. d, Fitting $M\propto L^{-a}$ for fixed $\tau/L$, where the value of $a$ does not converge as $\tau/L$ increases.
S1.5 Guiding principles for choosing the value of $\tau/L^{z}$ in simulations
In the main text, we adopt the form $\tau=\text{(const.)}\times L^{z}$ to identify quantum critical points. We use coefficients $0.3$ in Figs. 1bc and 2ab, and $0.25$ in Figs. 3bc. When choosing these specific coefficients, we carefully balanced three considerations: the severity of the sign problem, the validity of the scaling theory, and the numerical precision of the estimated critical point. In general, the following guiding principles apply:
  1. 1.
    The severity of the sign problem is the most important factor. Usually, slightly increasing the value of $\tau/L^{z}$ makes the sign problem significantly worse, and vice versa. For example, if the coefficient is increased from $0.3$ to $0.6$, the sign problem becomes notably more severe, significantly reducing the accessible system size. From this guiding principle, $\tau/L^{z}$ should be chosen as small as possible.
  2. 2.
    Another important factor is the validity of the scaling theory. If $\tau/L^{z}$ is too small, the nonequilibrium scaling theory may not hold well at currently accessible finite sizes. Specifically, for example, when it is taken below 0.1, from Figs. 1d, 2c, and 3d one can see that the data collapse gradually starts to break down, meaning finite-size effects become significant. This condition is relatively flexible. As long as $\tau/L^{z}$ is above a certain threshold, the scaling holds well, and variations in the coefficient do not significantly affect the results. Values such as $0.25$, $0.3$, and $0.5$ are typically used in the literature [79, 50], where estimates of critical points at different $\tau/L^{z}$ values were found to agree within statistical error, providing a useful self-consistency check.
  3. 3.
    Another somewhat subtle factor is the resolution of the critical point. Even within the valid scaling regime, choosing a smaller $\tau/L^{z}$ broadens the nonequilibrium critical region. This broadening manifests as a diminished dependence of the dimensionless correlation length ratio $R$ on system size $L$ over a wide range of the tuning parameter $U$ near $U_{c}$, making the crossing point of $R$ curves (such as in Fig. 1b) less sharp. This leads to reduced numerical precision in locating the critical point. Therefore, from this perspective, it is preferable to choose a somewhat larger $\tau/L^{z}$ to enhance resolution.
In summary, when applying the nonequilibrium approach to determine quantum critical points, one must balance these considerations within the valid scaling regime. The goal is to choose a value of $\tau/L^{z}$ that optimizes both precision and accuracy. In our experience, the severity of the sign problem is often the most decisive factor.
S2 More details for the single-Dirac-fermion Hubbard model
S2.1 Sign problem behaviors